OpenStudy (anonymous):

After blast-off, the sensors have picked up an incoming meteor. The Mathonauts frantically start preparing for impact. If the linear equation of the meteor’s path is known, describe to the captain how to solve where your equation from question 1 and the meteor’s path will cross. Explain any possible methods used in discovering a solution.

OpenStudy (anonymous):

I got told that I need 2-3 ways to find them. I know one is by using substition.

OpenStudy (perl):

use elimination

OpenStudy (anonymous):

Can you explain to me how that will help me find the path of the meteor please? thank you!

OpenStudy (jdoe0001):

ahemm what are the equations ?

OpenStudy (anonymous):

My equation is y=2x+1. The equation of the meteor's path is not given.

OpenStudy (jdoe0001):

elimination method is done by ADDING both equations vertically, once all variables are aligned let's say the meteor's path is y = -2x +25 so y = 2x + 1 y = -2x + 25 ------------ 2y = 0 + 26 and you can see right off what "y" would be, then you'd get "x" by substituting "y" on either equation

OpenStudy (jdoe0001):

by ADDING the equations vertically like so, you're in fact ELIMINATING one of the variables

OpenStudy (jdoe0001):

now, if you GRAPH both equations and the graphs touch each other or INTERSECT the solution is the (x, y) coordinate values at that INTERSECTION POINT that'd be the other way

OpenStudy (anonymous):

I see! So I can figure out the point of intersection of the meteor and the ship by using Substitution and Elimination. Thank you!

OpenStudy (jdoe0001):

yw